# Muon beams towards muonium physics: progress and prospects

### v1

@@ II.7 Muon Campus at FNAL
The proton accelerator system at the Muon Campus delivers an 8.89 GeV/ kinetic energy proton beam supplied by the Recycler ring to the AP0 target station initially [Ganguly:2022ufq]. The Muon Campus hosts both the Muon [Muong-2:2025xyk] and Mu2e [Atanov:2018ich] experiments, aiming at the discovery of BSM physics. For the Muon experiment, the 3.1 GeV muon beam is injected into the Delivery Ring (DR). After 4 turns in the DR, protons are kicked out to the beam dump, and the remaining muon beam is extracted to the Muon storage ring. For the Mu2e experiment, the proton beam passes the AP0 target and enters the DR directly. The DR will resonantly extract the proton beam to the Mu2e target, producing a PW muon beam. Additionally, several potential experiments at FNAL, including fixed-target muon experiments and muEDM studies, are also proposed, considering the Proton Improvement Plan-II (PIP-II) [Stanek:2023bae].
@@ II.8.1 HIMB at PSI
The High-Intensity Muon Beams (HIMB) will be constructed with the Isotope and Muon Production using Advanced Cyclotron and Target (IMPACT) project at PSI [impact2022]. Based on the original proton beam, the Target M station, including its beamlines M1 and M3 mentioned in Section II.2, will be replaced by a new Target H station with MUH2 and MUH3 beamlines as shown in Fig. 11. The goal of HIMB is to deliver the most intense CW surface muon beam at a flux of /s for both particle physics and material science experiments [Aiba:2021bxe]. This rate is expected to be achieved using a thicker target along with capture solenoids positioned near the target. The thickness of the new muon production target is optimized to be 20 mm, corresponding to effectively 100 mm along the beam direction. The capture solenoids are similar to the ones of the E4 beamline at PSI, and are expected to be resistant to radiation damage.
@@ II.8.2 MELODY at CSNS
The CSNS accelerator provides 1.6 GeV protons at a repetition rate of 25 Hz. Taking advantage of the CSNS-II upgrade plan, which increases the beam power from 100 kW to 500 kW, the construction of the Muon station for sciEnce, technoLOgy and inDustry (MELODY) has been proposed, with operation expected to begin by 2029 [Hong:2024gnp, Bao:2023nup]. The muon target station consists of two beamlines for surface muons and decay muons. Muons will be produced at the rate of /s using a 20 mm thick graphite target. However, the designed repetition frequency of 1 Hz is relatively low, and is expected to be increased to 5 Hz in the future upgrade plan [Chen:2023uvp].
@@ II.8.4 HIAF
II.8.5 RAMIS at RAON
@@ II.9 Summary of global muon beams
Over the past few decades, the development of high-power proton accelerators has provided opportunities for the construction of muon beam facilities. Taking advantage of high flux, high polarization, and controllable beam parameters, accelerator-based muon sources have become a fundamental platform for modern muon science and its applications, producing muons that cover an energy range from keV to GeV. Table 2 lists the global muon facilities that have been constructed, are under construction, or are being planned.
These facilities can be categorized into continuous wave (CW) and pulsed wave (PW) muon sources according to their time structures. In particular, the cyclotrons can deliver proton beams with high repetition rates. Due to the relatively long pion lifetime (26 ns), the emitted secondary muons are time-smeared, resulting in a CW beam. For CW beams, muons can be approximately regarded as entering the detector system and decaying individually. Therefore, under a good time resolution and an appropriate beam intensity, the “pile-up” effect can be effectively avoided. In the case of PW beams, muons arrive in bunches, each containing a large number of muons, which can significantly enhance the statistics of the experiment. Moreover, the time structure of the beam can provide an external trigger for event selection. However, “pile-up” effects may occur during the beam-on period.
| Mode | Facility | Name | Proton energy (GeV) | Proton power (MW) | Max rate level (s) | ||
@@
| PW | J-PARC/Japan | MUSE | 3/8 | 1 | |||
| ISIS/UK | EC | 0.8 | 0.18 | – | 10 | ||
@@
| FNAL/USA | Muon Campus | 8 | – | – | |||
| CSNS/China | MELODY | 1.6 | – | ||||
| ORNL-SNS/USA | – | 1.3 | 1.4/2.8 | ||||
| CW | PSI/Switerland | SS | 0.59 | 1.4 | |||
@@
| RNCP/Japan | MuSIC | 0.4 | – | ||||
| TRIUMF/Canada | CMMS | 0.52 | – | ||||
| CiADS/China | MuST | 0.6 | 0.3/3 | – | |||
| HIAF/China | – | 9.3 | – | ||||
| RAON/Korea | RAMIS | 0.6 | 0.1/0.4 | – | – | ||
Surface muons are the most widely used type in both particle physics and material science. Their main advantages include low energy, high polarization (), and monoenergetic. Slow muons can be produced by further moderating surface muons, and are applied in fields such as muon cooling, muonium physics, and surface physics studies. However, only a few beamlines are currently available for actual experiments. Related works mainly conducted in Switzerland, the UK, and Japan, but further optimization and testing are still needed. On the other hand, decay muons have high energy and stronger transmitting capabilities, making them suitable for muon scattering processes and muon tomography studies.
In the rare process searches and spectroscopic studies, the muon beam intensity is one of the most critical parameters. The reported beam intensities of existing and proposed muon facilities are summarized in Table 2. In future upgrades and construction plans for muon sources, several approaches are under consideration to enhance muon beam intensity, including increasing the proton beam power on target, optimizing the muon production targets, and improving the design of capture solenoids. The next-generation muon facilities, such as HIMB and MuST, are expected to increase the continuous surface muon intensity to the order of /s, which will significantly improve the precision of muon physics experiments.
@@ III Future prospects of muon beams
With the constant development of muon beams and the growing demands of experimental muon physics, future muon beams are expected to move toward more ambitious parameters: higher intensity, extreme high or low momentum region, and micro phase space [Nagamine:2003vx]. High-energy muon beams are also considered in the neutrino factories for an advanced source of muon neutrinos, providing a powerful tool to the current accelerator-based neutrino experiments such as T2K, NOvA, DUNE, and HyperK [Bogomilov:2014koa, T2K:2025wet, DUNE:2020jqi, Hyper-KamiokandeProto-:2015xww, An:2025lws]. The beamlines themselves are expected to be more cost-effective and more compact. Meanwhile, the study of advanced target materials could lead to a fundamental enhancement in muon yield and a reduction in positron contamination. For example, Ref. [Cai:2023caf] reports a liquid lithium target scheme that achieves a substantial increase in the muon rate. For the state-of-the-art muon collider prospect, significant challenges remain in the cooling and acceleration of muons. Here, we provide a brief review and discussion about several novel methods and techniques that are currently gaining attention.
@@ III.1.1 Higher energy ion or proton driven production
The proposed Neutrinos from Stored Muons (nuSTORM) facility, which plans to produce and store muon beams in a ring with central momenta between 1–6 GeV, taking advantage of the 100 GeV proton beam from the CERN Super Proton Synchrotron (SPS) [Ahdida:2020whw, nuSTORM:2025tph]. The storage ring scheme is also adopted by the Muon Campus at FNAL [Ganguly:2022ufq]. The M2 beamline at the CERN SPS delivers high energy muons with momenta up to 280 GeV [Bernhard:2019jqz]. This beamline hosts two fixed-target experiments: NA64, which searches for the invisible states via muon scattering processes [Crivelli:2025pjb], and MUonE, which aims to precisely measure the hadronic component of the muon anomaly [MUonE:2019qlm].
The mentioned facilities above can easily produce multi-GeV muons using the powerful heavy-ion or proton accelerators. However, due to the extended lifetime of pions, these facilities require large-scale beamline designs to ensure that pions can finally decay into muons. The contamination of muon beams could also be a possible challenge for both the beamline design and the experimental applications. The application of this type of muon source is therefore limited. Notably, a new approach has emerged in which muons are first “cooled” to significantly reduce their phase space, known as muon cooling. Then, the muons will be re-accelerated to high energies. This is now believed to be a critical step in the journey to future multi-TeV muon colliders.
@@ III.1.2 Accelerator-based electron driven production
In order to reduce the scale and cost of muon sources and to broaden their application scenarios, K. Nagamine et al. proposed a compact muon source design based on a 300 MeV, 10 A electron accelerators for SR application [Nagamine:2009zz]. When high-energy electrons interact with the target material nuclei, muons can be produced through the following processes [Blomqvist:1976mq, Nagamine:2009zz]:
@@
High-energy electron linac is vital for particle physics, material science, and biological science, especially as a key component of synchrotron radiation and X-ray free electron laser (XFEL) facilities [He_2014, Zhao_2017]. In recent years, there have been several feasibility studies for muon beamlines driven by high-energy electron beams through the mentioned processes. The currently under construction Shanghai High repetition rate XFEL and Extreme Light facility (SHINE) will deliver an 8 GeV electron beam operating at a 1 MHz repetition rate. A high-repetition-rate muon beam based on SHINE has been proposed, which is expected to generate /bunch using a thick tungsten target [Liu:2025ejy]. Similarly, the Continuous Electron Beam Accelerator Facility (CEBAF) at Jefferson Lab has proposed a baseline design of a beamdump facility, which would enable the construction of a 6 GeV muon beamline by exploiting its electron beam up to 12 GeV [Achenbach:2025ynn]. The muon rate is expected to reach /s level. For the direct annihilation mechanism, the Low EMittance Muon Accelerator (LEMMA) concept was proposed in Ref. [Boscolo:2020xfv]. In this scheme, 45 GeV positrons are used to reach the muon pair production threshold, and a thin liquid lithium target is employed. By simulation study, an emittance of m rad and a rate of /bunch could be obtained.
In summary, the muon production driven by an accelerator-based electron beam can not only help to reduce the overall size and cost of the facility but also suppress radiation contamination, including neutrons. However, such sources inherently suffer from a lower muon rate compared with proton-driven facilities due to the much smaller production cross sections. At present, no dedicated beamline of this type exists worldwide, and substantial development is still required.
@@ III.1.3 Laser-based electron driven production
The concept of wakefield acceleration was first introduced by T. Tajima and J.M. Dawson in 1979 [Tajima:1979bn]. They proposed that the ponderomotive force of an intense laser pulse could drive a plasma wake capable of accelerating trapped electrons to high energies over extremely short distances—a process often described as “surfing”. Limited by the laser technology available at that time, their work was only based on theoretical calculation and simulation. Substantial experimental progress was achieved in 2004, when three independent groups reported the generation of high-quality, quasi-monoenergetic electron beams with energies around 100 MeV, marking the first clear demonstration of LWFA [Mangles:2004ta, Geddes:2004tb, Faure:2004tc]. In 2024, A. Picksley et al. reported the acceleration of quasi-monoenergetic electrons up to 9.2 GeV using a 500 TW laser at the BErkeley Laboratory Laser Accelerator (BELLA) facility, representing the highest energy of LWFA electrons reported to date [Picksley:2024cdd]. As petawatt-class lasers and staged acceleration techniques continue to mature, further increases in beam energy are expected.
High-energy electrons produced via wakefield acceleration can be used for muon generation, and several simulation studies have previously explored this possibility, as reported in Ref. [Rao:2018njj, Calvin_2023]. In 2025, F. Zhang et al. employed the 1-PW laser of the Shanghai Superintense Ultrafast Laser Facility (SULF) to accelerate electrons to 0.4–1.5 GeV for muon production [Zhang:2024axy]. The electrons were then transported to a 12 cm-thick lead target, behind which two liquid scintillators shielded by lead were installed. By measuring the time difference between the fast and slow signals of muons and their Michel electrons, then fitting the average muon lifetime, the observation of muon production was confirmed. Geant4 simulations further estimated a muon yield of 0.01 muons per incident electron, representing the first proof-of-principle demonstration of a laser-driven muon source. Subsequently, a research group at LBNL reported a similar experiment at BELLA and observed muons downstream of a 2.1 cm tungsten target as Fig. 12 [Terzani:2024neq]. Furthermore, at the Extreme Light Infrastructure-Nuclear Physics (ELI-NP) facility, experiments were conducted using a 10 cm lead target, in which two dipole magnets transported the produced muons to the detector region [Calvin:2025huk]. Differing from the previous two studies, this experiment uses silicon pixel sensors read out by Timepix3 instead of scintillators, making it possible to identify muons through reconstructed track length and energy deposition.

### v2

@@ II.7 Muon Campus at FNAL
The proton accelerator system at the Muon Campus (as shown in Fig.˜10) delivers an 8.89 GeV/ proton beam supplied by the Recycler ring to the AP0 target station initially [Ganguly:2022ufq]. The Muon Campus hosts both the Muon [Muong-2:2025xyk] and Mu2e [Atanov:2018ich] experiments, aiming at the discovery of BSM physics. For the Muon experiment, the 3.1 GeV muon beam is injected into the Delivery Ring (DR). After 4 turns in the DR, protons are kicked out to the beam dump, and the remaining muon beam is extracted to the Muon storage ring. For the Mu2e experiment, the proton beam passes the AP0 target and enters the DR directly. The DR will resonantly extract the proton beam to the Mu2e target, producing a PW muon beam. Additionally, several potential experiments at FNAL, including fixed-target muon experiments and muon EDM studies, are also proposed, considering the Proton Improvement Plan-II (PIP-II) [Stanek:2023bae].
@@ II.8.1 HIMB at PSI
The High-Intensity Muon Beams (HIMB) will be constructed with the Isotope and Muon Production using Advanced Cyclotron and Target (IMPACT) project at PSI [impact2022]. Based on the original proton beam, the Target M station, including its beamlines M1 and M3 mentioned in Section˜II.2, will be replaced by a new Target H station with MUH2 and MUH3 beamlines as shown in Fig.˜11. The goal of HIMB is to deliver the most intense CW surface muon beam at a flux of /s for both particle physics and materials science experiments [Aiba:2021bxe]. This rate is expected to be achieved using a thicker target along with capture solenoids positioned near the target. The thickness of the new muon production target is optimized to be 20 mm, corresponding to effectively 100 mm along the beam direction. The capture solenoids are similar to the ones of the E4 beamline at PSI, and are expected to be resistant to radiation damage.
@@ II.8.2 MELODY at CSNS
The CSNS accelerator provides 1.6 GeV protons at a repetition rate of 25 Hz. Taking advantage of the CSNS-II upgrade plan, which increases the beam power from 100 kW to 500 kW, the construction of the Muon station for sciEnce, technoLOgy and inDustry (MELODY) has been proposed, with operation expected to begin by 2029 [Hong:2024gnp, Bao:2023nup]. The muon target station consists of two beamlines for surface muons and decay muons. Muons will be produced at the rate of /s using a copper target with dimensions of to improve heat dissipation performance [Qiang:2024knc]. However, the designed repetition frequency of 1 Hz is relatively low, and is expected to be increased to 5 Hz in the future upgrade plan [Chen:2023uvp].
@@ II.8.4 HIAF
II.8.5 HEF at J-PARC
In addition to MUSE, J-PARC is preparing a dedicated muon beamline for the COMET experiment, which searches for conversion [COMET:2018auw]. This beamline is located at the Hadron Experimental Facility (HEF), which operates with a 30 GeV proton beam extracted from the Main Ring accelerator at a maximum power of 84 kW [Takahashi:2025wqw]. The COMET experiment requires low-energy negative muons to be stopped in a target to form muonic atoms. Therefore, an 8 GeV operation mode of the proton beam will be used. The “bunched slow extraction” technique is employed to maintain a well-defined time structure, with a pulse width of about 100 ns and a separation greater than 1.17 s. For muon production and transport, COMET introduces an superconducting solenoid system demonstrated at the MuSIC facility (described in Section˜II.4) to obtain an intense low-momentum negative muon beam. Muons with momenta of about 40 MeV/ will be delivered to the detector region and stopped at the center of the aluminum target. The experiment will be carried out in two stages, Phase-I and Phase-II, requiring beam powers of 3.2 kW and 56 kW, respectively, and is expected to achieve a maximum muon intensity of up to /s.
II.8.6 RAMIS at RAON
@@ II.9 Summary of global muon beams
Over the past few decades, the development of high-power proton accelerators has provided opportunities for the construction of muon beam facilities. Taking advantage of high flux, high polarization, and controllable beam parameters, accelerator-based muon sources have become a fundamental platform for modern muon science and its applications, producing muons that cover an energy range from keV to GeV. Table˜2 lists the global muon facilities that have been constructed, are under construction, or are being planned.
These facilities can be categorized into continuous-wave (CW) and pulsed-wave (PW) muon sources according to their time structures. In particular, the cyclotrons can deliver proton beams with high repetition rates. Due to the relatively long pion lifetime (26 ns), the emitted secondary muons are time-smeared, resulting in a CW beam. For CW beams, muons can be approximately regarded as entering the detector system and decaying individually. Therefore, under a good time resolution and an appropriate beam intensity, the “pile-up” effect can be effectively avoided. However, some degree of “pile-up” effect will also occur when the muon intensity becomes too large (). In the case of PW beams, muons arrive in bunches, each containing a large number of muons, which can significantly enhance the statistics of the experiment. Moreover, the time structure of the beam can provide an external trigger for event selection. However, “pile-up" effects may occur during the beam-on period.
| Mode | Facility | Name | Proton energy (GeV) | Proton power (kW) | Max rate level (s) | ||
@@
| PW | J-PARC/Japan | MUSE | 3/8 | 1000 | |||
| HEF | 8 | 3.2/56 | |||||
| ISIS/UK | EC | 0.8 | 180 | – | 10 | ||
@@
| FNAL/USA | Muon Campus | 8 | 25 | – | – | ||
| CSNS/China | MELODY | 1.6 | 20 | – | |||
| HIAF-HIRIBL/China | – | 9.3 | – | – | – | ||
| ORNL-SNS/USA | – | 1.3 | 1400/2800 | ||||
| CW | PSI/Switerland | SS | 0.59 | 1400 | |||
@@
| RNCP/Japan | MuSIC | 0.4 | 0.4 | – | |||
| TRIUMF/Canada | CMMS | 0.52 | 75 | – | |||
| CiADS/China | MuST | 0.6 | 300/3000 | – | |||
| RAON/Korea | RAMIS | 0.6 | 100/400 | – | – | ||
Surface muons are the most widely used type in both particle physics and materials science. Their main advantages include low energy, high polarization (), and monoenergetic. Slow muons can be produced by further moderating surface muons, and are applied in fields such as muon cooling, muonium physics, and surface physics studies. However, only a few beamlines are currently available for actual experiments. Related works mainly conducted in Switzerland, the UK, and Japan, but further optimization and testing are still needed. On the other hand, decay muons have high energy and stronger transmitting capabilities, making them suitable for muon scattering processes and muon tomography studies.
In the rare process searches and spectroscopic studies, the muon beam intensity is one of the most critical parameters. The reported beam intensities of existing and proposed muon facilities are summarized in Table˜2. In future upgrades and construction plans for muon sources, several approaches are under consideration to enhance muon beam intensity, including increasing the proton beam power on target, optimizing the muon production targets, and improving the design of capture solenoids. The next-generation muon facilities, such as HIMB and MuST, are expected to increase the continuous surface muon intensity to the order of /s, which will significantly improve the precision of muon physics experiments.
@@ III Future prospects of muon beams
With the constant development of muon beams and the growing demands of experimental muon physics, future muon beams are expected to move toward more ambitious parameters: higher intensity, extreme high or low momentum region, micro phase space, and beyond [Nagamine:2003vx, Cywinski:2009zz]. High-energy muon beams are also considered in the neutrino factories for an advanced source of muon neutrinos, providing a powerful tool to the current accelerator-based neutrino experiments such as T2K, NOvA, DUNE, and HyperK [Bogomilov:2014koa, T2K:2025wet, DUNE:2020jqi, Hyper-KamiokandeProto-:2015xww, An:2025lws]. The beamlines themselves are expected to be more cost-effective and more compact. Meanwhile, the study of advanced target materials could lead to a fundamental enhancement in muon yield and a reduction in positron contamination. For example, Ref. [Cai:2023caf] reports a liquid lithium target scheme that achieves a substantial increase in the muon rate. For the state-of-the-art muon collider prospect, significant challenges remain in the cooling and acceleration of muons. Here, we provide a brief review and discussion of several novel methods and techniques that are currently gaining attention.
@@ III.1.1 Higher energy ion or proton driven production
The proposed Neutrinos from Stored Muons (nuSTORM) facility plans to produce and store muon beams in a ring with central momenta between 1–6 GeV, taking advantage of the 100 GeV proton beam from the CERN Super Proton Synchrotron (SPS) [Ahdida:2020whw, nuSTORM:2025tph]. The storage ring scheme is also adopted by the Muon Campus at FNAL [Ganguly:2022ufq]. The M2 beamline at the CERN SPS delivers high energy muons with momenta up to 280 GeV [Bernhard:2019jqz]. This beamline hosts two fixed-target experiments: NA64, which searches for the invisible states via muon scattering processes [Crivelli:2025pjb], and MUonE, which aims to precisely measure the hadronic component of the muon anomaly [MUonE:2019qlm].
The mentioned facilities above can easily produce multi-GeV muons using the powerful heavy-ion or proton accelerators. However, due to the extended lifetime of pions, these facilities require large-scale beamline designs to ensure that pions can finally decay into muons. The contamination of muon beams could also be a possible challenge for both the beamline design and the experimental applications. The application of this type of muon source is therefore limited. Notably, a new approach has emerged in which muons are first “cooled" to significantly reduce their phase space, known as muon cooling. Then, the muons will be re-accelerated to high energies. This is now believed to be a critical step in the journey to future multi-TeV muon colliders.
@@ III.1.2 Accelerator-based electron driven production
In order to reduce the scale and cost of muon sources and to broaden their application scenarios, K. Nagamine et al. proposed a compact muon source design based on a 300 MeV, 10 A electron accelerator for SR applications [Nagamine:2009zz]. When high-energy electrons interact with the target material nuclei, muons can be produced through the following processes [Blomqvist:1976mq, Nagamine:2009zz]:
@@
A high-energy electron linac is vital for particle physics, materials science, and biological science, especially as a key component of synchrotron radiation and X-ray free electron laser (XFEL) facilities [He_2014, Zhao_2017]. In recent years, there have been several feasibility studies for muon beamlines driven by high-energy electron beams through the mentioned processes. The Shanghai High repetition rate XFEL and Extreme Light facility (SHINE) currently under construction will deliver an 8 GeV electron beam operating at a 1 MHz repetition rate. A high-repetition-rate muon beam based on SHINE has been proposed, which is expected to generate /bunch using a thick tungsten target [Liu:2025ejy]. Similarly, the Continuous Electron Beam Accelerator Facility (CEBAF) at Jefferson Lab has proposed a baseline design of a beamdump facility, which would enable the construction of a 6 GeV muon beamline by exploiting its electron beam up to 12 GeV [Achenbach:2025ynn]. The muon rate is expected to reach /s level. For the direct annihilation mechanism, the Low EMittance Muon Accelerator (LEMMA) concept was proposed in Ref. [Boscolo:2020xfv]. In this scheme, 45 GeV positrons are used to reach the muon pair production threshold, and a thin liquid lithium target is employed. By simulation study, an emittance of m rad and a rate of /bunch could be obtained.
In summary, the muon production driven by an accelerator-based electron beam can not only help reduce the overall size and cost of the facility but also suppress radiation contamination, including neutrons. However, such sources inherently suffer from a lower muon rate compared with proton-driven facilities due to the much smaller production cross sections. At present, no dedicated beamline of this type exists worldwide, and substantial development is still required.
@@ III.1.3 Laser-based electron driven production
The concept of wakefield acceleration was first introduced by T. Tajima and J.M. Dawson in 1979 [Tajima:1979bn]. They proposed that the ponderomotive force of an intense laser pulse could drive a plasma wake capable of accelerating trapped electrons to high energies over extremely short distances—a process often described as “surfing". Limited by the laser technology available at that time, their work was only based on theoretical calculation and simulation. Substantial experimental progress was achieved in 2004, when three independent groups reported the generation of high-quality, quasi-monoenergetic electron beams with energies around 100 MeV, marking the first clear demonstration of LWFA [Mangles:2004ta, Geddes:2004tb, Faure:2004tc]. In 2024, A. Picksley et al. reported the acceleration of quasi-monoenergetic electrons up to 9.2 GeV using a 500 TW laser at the BErkeley Laboratory Laser Accelerator (BELLA) facility, representing the highest energy of LWFA electrons reported to date [Picksley:2024cdd]. As petawatt-class lasers and staged acceleration techniques continue to mature, further increases in beam energy are expected.
High-energy electrons produced via wakefield acceleration can be used for muon generation, and several simulation studies have previously explored this possibility, as reported in Ref. [Rao:2018njj, Calvin_2023]. In 2025, F. Zhang et al. employed the 1-PW laser of the Shanghai Superintense Ultrafast Laser Facility (SULF) to accelerate electrons to 0.4–1.5 GeV for muon production [Zhang:2024axy]. The electrons were then transported to a 12 cm-thick lead target, behind which two liquid scintillators shielded by lead were installed. By measuring the time difference between the fast and slow signals of muons and their Michel electrons, then fitting the average muon lifetime, the observation of muon production was confirmed. Geant4 simulations further estimated a muon yield of 0.01 muons per incident electron, representing the first proof-of-principle demonstration of a laser-driven muon source. Subsequently, a research group at LBNL reported a similar experiment at BELLA and observed muons downstream of a 2.1 cm tungsten target as Fig.˜12 [Terzani:2024neq]. Furthermore, at the Extreme Light Infrastructure-Nuclear Physics (ELI-NP) facility, experiments were conducted using a 10 cm lead target, in which two dipole magnets transported the produced muons to the detector region [Calvin:2025huk]. Differing from the previous two studies, this experiment uses silicon pixel sensors read out by Timepix3 instead of scintillators, making it possible to identify muons through reconstructed track length and energy deposition.